OpenStudy (jokersmart):

What is 4.62 expressed as the quotient of two integers in simplest form?

OpenStudy (jokersmart):

\[\frac{ 231 }{ 50 } \] right?

OpenStudy (mathstudent55):

Just from looking at the number, you should be able to tell that it is equal to \(\dfrac{462}{100} \) Then you reduced correctly to \(\dfrac{231}{50} \) which is fully reduced. You are correct.

OpenStudy (jokersmart):

Thank you. (: